# Deep FlexQP: Accelerated Nonlinear Programming via Deep Unfolding

### v1

@@ 4.2 Supervised Learning
For training Deep OSQP variants, we adopt the supervised learning approach from saravanos2025deep. Let $x^{k}(\theta)$ be the $k^{\text{th}}$ iterate of Deep OSQP parameterized by $\theta$. The training objective is the weighted sum of the optimality gaps between the iterates $x^{k}(\theta)$ and the optimal solution $x^{*}$:
@@
where $\xi=(x,y_{I},y_{E})$. By including the Lagrange multipliers here, we are able to enforce the Deep FlexQP optimizer to select penalty parameters that meet the conditions of Theorem 3.1, namely that $\mu_{I}\geq\left\lVert y_{I}^{*}\right\rVert_{\infty}$ and $\mu_{E}\geq\left\lVert y_{E}^{*}\right\rVert_{\infty}$. This is due to the fact that the Lagrange multipliers of Deep FlexQP $y_{I}(\theta)$ and $y_{E}(\theta)$ are upper-bounded (in absolute value) by the current selection of $\mu$ (see Theorem 3.3). An ablation studying the effect of this loss is provided in Appendix N.
@@ 4.3 PAC-Bayes Generalization Bounds
Recent approaches have been proposed for establishing generalization bounds for guaranteeing the performance of learning-to-optimize methods, including a binary loss approach from sambharya2025data as well as a more informative progress metric by saravanos2025deep, given as
@@
These approaches can be used to construct PAC-Bayes generalization bounds on the mean performance of the optimizer that hold with high probability. Nevertheless, a limitation of the resulting PAC-Bayes bound from Equation 13 is that it assumes that the losses can fall anywhere within in the range $[0,1]$, despite the fact that, in practice, most of the final optimality gaps fall very close to 0 (on the order of 1e2110-2 and smaller). In other words, the loss in Equation 13 does not properly account for the scale of the errors, and as a result, obtaining a meaningful bound might require exponentially more training samples. For example, Figure 4(a) shows that training for this generalization bound loss results in a bound that is uninformative since it sits above even the vanilla optimizers and does not capture the behavior well at small errors.
@@
To address this issue, we design a loss that is zero when the learner performs as well as or better than the optimal solution, and increases linearly as the performance decreases on a log-scale; see Figure 3 for a visualization. Furthermore, we penalize distance from the optimal solution with respect to the norm of the QP residuals, which better takes into account the problem scale:
@@
where $R(\xi)\vcentcolon=(Px+q+G^{\top}y_{I}+A^{\top}y_{E},\max(Gx-h,0),Ax-b)$ computes the residuals of the original QP in Equation 3. As intended, training for this loss better captures the performance when the residuals are very small. Results are presented in Figure 4(b) and Appendix L.
(a) Deep FlexQP trained for the generalization bound loss Equation 13.
(b) Deep FlexQP trained for our proposed generalization bound loss Equation 14.
Figure 4: Optimizer comparison on 1000 test LASSO problems. Training using our log-normalized loss Equation 14 results in a substantially more informative performance guarantee.
@@ 5.1 Small- to Medium-Scale QPs
We apply our deep-unfolded methodology to a benchmark suite of QPs including portfolio optimization problems from finance, classification and regression problems from machine learning, and linear optimal control problems. The results are presented in Figure 5, and details on the problem representations as well as the data generation processes are provided in Appendix G. All experiments were run using PyTorch on a system with an Intel i9-13900K processor and an NVIDIA RTX 4090 GPU. In the following plots, OSQP and FlexQP are the best performing versions of OSQP and FlexQP, respectively, found using a hyperparameter search (details in Appendix I). Deep OSQP is the approach from saravanos2025deep, Deep OSQP — RLQP Parameterization is the parameterization from ichnowski2021accelerating, and Deep OSQP — Improved is the best-of-both-worlds version of deep-unfolded OSQP described in Section 4. Finally, Deep FlexQP is our proposed deep-unfolded FlexQP optimizer with LSTM policy parameterization and trained using the loss Equation 12. Each model is trained for 500 epochs on 500 problems (except for the random QP classes, which use 2000 problems) and evaluated on 1000 test problems. We also perform an extensive timing comparison and analysis between all of the above optimizers, presented in Appendix I.
Figure 5: Performance comparison of learned deep optimizers and their non-learned counterparts on small- to medium-scale QPs. Our improved version of Deep OSQP outperforms the baselines, while Deep FlexQP consistently surpasses the rest of the methods in terms of convergence to the optimal QP solution.
@@ 5.2 Large-Scale QPs
[…]
@@ 5.3 Nonconvex Nonlinear Programming using SQP
Lastly, we apply Deep FlexQP as a submodule in SQP to solve nonconvex NLPs arising from nonlinear optimal control and nonlinear predictive safety filter problems. Training for the generalization bound loss proposed in Equation 14 yields a numerical certificate of performance that we use when designing the SQP method. The results are provided in Figure 1 and Figure 7. Further details are provided in Appendix H.
@@ 6 Conclusion
We present FlexQP, a flexible QP solver that can solve any convex QP with minimal assumptions on the constraints. FlexQP always returns a solution that minimizes the constraint violation, and thus can be used as a robust QP solver in SQP. Our accelerated variant, Deep FlexQP, outperforms other traditional optimizers and learned approaches in terms of convergence, both in number of iterations and solve time. Using Deep FlexQP as a submodule in SQP provides a substantial speedup over traditional approaches, while also allowing for a graceful recovery when an infeasible QP subproblem is encountered. Some potential future extensions include learning warm-starts for our solver and applying it to the distributed QP setting explored in saravanos2025deep.
Appendix A Proof of Theorem 3.1
@@
Lemma A.1.
The relaxed QP Equation 6 can equivalently be expressed as the following optimization:
$\min_{x}\quad\phi(x;\mu_{I},\mu_{E})\vcentcolon=\frac{1}{2}x^{\top}Px+q^{\top}x+\mu_{I}\left\lVert(Gx-h)_{+}\right\rVert_{1}+\mu_{E}\left\lVert Ax-b\right\rVert_{1}.$ (15)
Proof Convert the optimization to the form without slack variables. This can be equivalently derived by directly relaxing the constraints of the original QP Equation 3 using $\ell_{1}$ penalties.
@@
    if $x^{*}$ solves Equation 3, then $x^{*}$ solves Equation 15, and
@@
    if $\hat{x}$ solves Equation 15 then $\hat{x}$ solves Equation 3.
@@
Let $x^{*}$ solve Equation 3. For any $x\in\mathbb{R}^{n}$ we have that
@@
The step from Equation 16b to Equation 16c follows from the fact that $\mu_{I}\geq\left\lVert y_{I}^{*}\right\rVert_{\infty}$ and $\mu_{E}\geq\left\lVert y_{E}^{*}\right\rVert_{\infty}$, and Equation 16d follows from the definition of the infinity norm. Equation 16e follows from the definition of ReLU and the absolute value and Equation 16f is implied by the linearity of the constraints. Obtaining Equation 16g follows from the complementary slackness condition $y_{I,i}^{*}(g_{i}^{\top}x_{i}^{*}-h_{i})=0$ for all $i=1,\ldots,m$, and feasibility $a_{i}^{\top}x_{i}^{*}-b_{i}=0$ for all $i=1,\ldots,p$. Equation 16i comes from the stationary condition $Px^{*}+q+G^{\top}y_{I}^{*}+A^{\top}y_{E}^{*}=0$. Finally, obtaining Equation 16k follows from the positive semidefiniteness of $P$.
Thus, we have shown that $\phi(x^{*};\mu_{I},\mu_{E})\leq\phi(x;\mu_{I},\mu_{E})$ for any $x$, which implies that $x^{*}$ minimizes $\phi(x;\mu_{I},\mu_{E})$ and therefore solves Equation 15.

### v2

@@ 4.2 Supervised Learning
For training Deep OSQP variants, we adopt the supervised learning approach from Saravanos et al. (2025). Let $x^{k}(\theta)$ be the $k^{\text{th}}$ iterate of Deep OSQP parameterized by $\theta$. The training objective is the weighted sum of the optimality gaps between the iterates $x^{k}(\theta)$ and the optimal solution $x^{*}$:
@@
where $\xi=(x,y_{I},y_{E})$. By including the Lagrange multipliers here, we are able to enforce the Deep FlexQP optimizer to select penalty parameters that meet the conditions of Theorem 3.1, namely that $\mu_{I}\geq\left\lVert y_{I}^{*}\right\rVert_{\infty}$ and $\mu_{E}\geq\left\lVert y_{E}^{*}\right\rVert_{\infty}$. This is due to the fact that the Lagrange multipliers of Deep FlexQP $y_{I}(\theta)$ and $y_{E}(\theta)$ are upper-bounded (in absolute value) by the current selection of $\mu$ (see Theorem 3.3). An ablation studying the effect of this loss is provided in Appendix O.
@@ 4.3 PAC-Bayes Generalization Bounds
Recent approaches have been proposed for establishing generalization bounds for guaranteeing the performance of learning-to-optimize methods, including a binary loss approach from Sambharya and Stellato (2025) as well as a more informative progress metric by Saravanos et al. (2025), given as
@@
These approaches can be used to construct PAC-Bayes generalization bounds on the mean performance of the optimizer that hold with high probability. Nevertheless, a limitation of the resulting PAC-Bayes bound from Eq. 13 is that it assumes that the losses can fall anywhere within the range $[0,1]$, despite the fact that, in practice, most of the final optimality gaps fall very close to 0 (on the order of $10^{-2}$ and smaller). In other words, the loss in Eq. 13 does not properly account for the scale of the errors, and as a result, obtaining a meaningful bound might require exponentially more training samples. For example, Fig. 4(a) shows that training for this generalization bound loss results in a bound that is uninformative since it sits above even the vanilla optimizers and does not capture the behavior well at small errors.
To address this, we design a loss that is zero when the learner performs as well as or better than the optimal solution, and increases linearly as the performance decreases on a log-scale; see Fig. 3 for a visualization. Furthermore, we penalize distance from the optimal solution with respect to the norm of the QP residuals, which better takes into account the problem scale:
@@
where $R(\xi)\vcentcolon=(Px+q+G^{\top}y_{I}+A^{\top}y_{E},\max(Gx-h,0),Ax-b)$ computes the residuals of the original QP in Eq. 3. As intended, training for this loss better captures the performance when the residuals are very small. Results are presented in Fig. 4(b) and Appendix M.
(a) Deep FlexQP trained for the generalization bound loss Eq. 13.
(b) Deep FlexQP trained for our proposed generalization bound loss Eq. 14.
Figure 4: Optimizer comparison on 1000 test LASSO problems. Training using our log-normalized loss Eq. 14 results in a substantially more informative performance guarantee.
@@ 5 Experimental Results
We evaluate our approach on three classes of problems of increasing complexity: small- to medium-scale QPs, large-scale QPs, and nonconvex nonlinear programs solved via SQP. All algorithms were implemented using dense batched matrices in PyTorch, and all experiments were run on a system with an NVIDIA RTX 4090 GPU. Across all experiments, we use a learning rate of $10^{-3}$ and a batch size of 50 (except for the large-scale problems, which use a batch size of 1, as larger batches exceeded the available GPU memory).
@@ 5.1 Small- to Medium-Scale QPs
We apply our deep-unfolding methodology to a benchmark suite of QPs including portfolio optimization problems from finance, classification and regression problems from machine learning, and linear optimal control problems. The results are presented in Fig. 5, and details on the problem representations as well as the data generation processes are provided in Appendix H. In the following plots, OSQP and FlexQP are the best performing versions of OSQP and FlexQP, respectively, found using a hyperparameter search (details in Appendix J). Deep OSQP is the approach from Saravanos et al. (2025), Deep OSQP — RLQP Parameterization is the parameterization from Ichnowski et al. (2021), and Deep OSQP — Improved is the best-of-both-worlds version of deep-unfolded OSQP described in Section 4. Finally, Deep FlexQP is our proposed deep-unfolded FlexQP optimizer with LSTM policy parameterization and trained using the loss Eq. 12. Each model is trained for 500 epochs on 500 problems (except for the random QP classes, which use 2000 problems) and evaluated on 1000 test problems. We also perform an extensive timing comparison and analysis between all of the above optimizers, presented in Appendix J.
Figure 5: Performance comparison of learned deep optimizers and their non-learned counterparts. Our improved version of Deep OSQP outperforms the baselines, while Deep FlexQP consistently surpasses the rest of the methods in terms of convergence to the optimal QP solution.
@@ 5.2 Large-Scale QPs
Next, we verify how well our methodology generalizes to large-scale QPs. In order to amortize the cost of training on these large-scale problems, we adopt a fine-tuning approach where the models trained on the small- to medium-scale problems from Section 5.1 are fine-tuned on a limited number of large-scale problems for a few epochs. Each model was fine-tuned on 100 training problems for 5 epochs. As each epoch takes roughly 3 hours to run, we estimate that training on the same number of problems and for the same number of epochs as the problems considered in Section 5.1 would take over 300 days, showing a clear benefit of the proposed fine-tuning approach.
Fig. 6 shows a comparison on 100 portfolio optimization and support vector machine (SVM) test problems with 10k variables and 10-20k constraints. Each optimizer is run until the infinity norm of the residuals falls below an absolute tolerance of $\varepsilon=10^{-3}$, with a timeout of 10 minutes. Optimizers use the indirect method to solve their first block ADMM update (see Appendix C). For the portfolio optimization problems, we report the average number of iterations each optimizer took to converge, and, for the SVM problems, we report the average number of conjugate gradient (CG) iterations that were necessary to solve the linear systems to a tolerance of $\varepsilon_{\text{CG}}=10^{-2}\cdot\varepsilon$, with full results provided in Appendix K. We observe that the fine-tuned Deep FlexQP outperforms all of the other optimizers. Surprisingly, the fine-tuning approach does not seem to work as well for the Deep OSQP variants. Notably, we tried fine-tuning each of the methods using either the optimality gap loss Eq. 11 or the Lagrange multiplier loss Eq. 12. While the two losses yielded comparable performance on the portfolio optimization problems, the Lagrange multiplier loss performed significantly worse on the SVM problems for the Deep OSQP variants. As such, the reported Deep OSQP results use the models trained with the optimality gap loss Eq. 11. These results suggest that the superior fine-tuning performance of Deep FlexQP is attributable to the FlexQP architecture itself — specifically, the elastic relaxation and LSTM parameterization — rather than the choice of loss function.
@@ 5.3 Nonconvex Nonlinear Programming using SQP
Lastly, we apply Deep FlexQP as a submodule in SQP to solve nonconvex NLPs arising from nonlinear optimal control and nonlinear predictive safety filter problems. Training for the generalization bound loss Eq. 14 yields a numerical certificate of performance that we use when designing the SQP method. Results are shown in Fig. 1 and Fig. 7. Further details are provided in Appendix I.
@@ 6 Conclusion
We present FlexQP, a flexible QP solver that natively handles infeasible subproblems by minimizing constraint violation when no feasible solution exists. This property makes FlexQP particularly well-suited as a submodule in SQP, where infeasible QP subproblems are common and typically require ad hoc recovery strategies. Our accelerated variant, Deep FlexQP, outperforms both traditional solvers and learned approaches, converging in fewer iterations and less time per solve. Generalization bounds provide a numerical certificate of performance, and we use these bounds to design SQP solvers for nonlinear optimal control and predictive safety filters. In our SQP experiments, Deep FlexQP provides a substantial speedup over traditional approaches while also enabling graceful recovery from infeasibility without additional machinery. Promising future directions include learning warm-starts for FlexQP and extending it to the distributed QP setting of Saravanos et al. (2025).
Acknowledgments
This work was supported by the Army Research Office Award #W911NF2010151. Alex Oshin was additionally supported by the Georgia Tech Aerospace Engineering Graduate Student Fellowship, and Augustinos Saravanos by the A. Onassis Foundation Scholarship.
Appendix A Extended Related Work
Quadratic Programming. Active-set methods, developed as extensions of the simplex method for QPs (Wolfe, 1959), can be efficiently warm-started but suffer from worst-case exponential complexity (Klee and Minty, 1970). Interior-point methods rose to prominence with the polynomial-time algorithms of Karmarkar (1984) and their extension to general convex programs via self-concordant barriers (Nesterov and Nemirovskii, 1994), replacing active-set methods as the dominant approach for small- to medium-scale problems (Wright, 2004). However, interior-point methods often do not scale well to high-dimensional problems and are difficult to warm-start. This has led to a growing interest in first-order methods in recent years. These methods are largely based on operator splitting methods such as ADMM, which is equivalent to Douglas-Rachford splitting applied to the dual (Gabay, 1983). First-order methods are appealing as they scale favorably to millions of variables and can be effectively warm-started, unlike interior-point methods. In addition, they are especially well-suited to applications such as MPC, where only moderate accuracy is required and the slow tail convergence of ADMM (He and Yuan, 2012) is less problematic.
SCS (O’Donoghue et al., 2016) is a first-order solver for general conic programs based on a homogeneous self-dual embedding. However, using this embedding is inefficient for QPs since it requires reformulating the quadratic cost as a second-order cone constraint. OSQP (Stellato et al., 2020) and COSMO (Garstka et al., 2021) are ADMM-based solvers specifically designed for QPs and general conic programs, respectively. COSMO can solve QPs as a special case, with the resulting algorithm being nearly identical to OSQP. We discuss these methods further below in the context of infeasibility detection.
The proposed FlexQP algorithm is a first-order method based on ADMM that has the same per-iteration complexity as OSQP and COSMO (see Appendix C), but uniquely handles infeasible QPs through an exact $\ell_{1}$ relaxation of the constraints (Section 3).
Penalty Methods and Constrained Optimization. Penalty methods have a long history in constrained nonlinear programming, starting from the early works by Courant (1943) proposing the quadratic penalty method, to the log and inverse barrier methods of Frisch (1955) and Carroll (1961). Fiacco and McCormick (1968) provided a unifying perspective on exterior-point penalty methods and interior-point barrier methods, and also began the analysis on primal-dual methods.
However, both exterior-point and interior-point methods suffer from poor numerical conditioning as the algorithm converges because the penalty parameter needs to asymptotically approach infinity (for exterior-point methods) or zero (for interior-point methods). This led to the development of exact penalty methods, based primarily on $\ell_{1}$ penalty functions (Eremin, 1966; Zangwill, 1967). These methods are more numerically stable because the penalty parameter need only be larger than the largest Lagrange multiplier. The necessary and sufficient conditions for exactness were given by Pietrzykowski (1969). The $\ell_{1}$ penalty function was also used by Han (1977) to prove the global convergence of SQP, foreshadowing the key role that exact penalties play in handling infeasible subproblems, as we discuss below. A modern discussion is given by Nocedal and Wright (2006) and we use these results to prove the exactness of our relaxed problem formulation (Theorem 3.1).
While exact penalty methods fixed the numerical ill-conditioning of previous approaches, the $\ell_{1}$ penalty function is non-differentiable at the boundary of the feasible region, resulting in a phenomenon known as the Maratos effect (Maratos, 1978). This sparked interest in methods that modify the SQP algorithm to preserve convergence, such as watchdog procedures (Chamberlain et al., 1982) and second-order corrections (Conn and Pietrzykowski, 1977; Fletcher, 1982).
Operator Splitting and ADMM. Concurrently, the augmented Lagrangian (AL) method, also known as the method of multipliers, was developed by Hestenes (1969) and Powell (1969). The alternating direction method of multipliers (ADMM) was subsequently developed by Glowinski and Marroco (1975) and Gabay and Mercier (1976). The popularity of ADMM increased significantly in the 2010s after its suitability for solving large-scale optimization problems was highlighted by Boyd et al. (2011). Since then, highly scalable ADMM-based methods have been developed for signal processing (Mateos et al., 2010; Zhang and Kwok, 2014), multi-agent robotics (Saravanos et al., 2021; 2023; Abdul et al., 2025), and power systems (Erseghe, 2014; Mhanna et al., 2018; Xu et al., 2018), among other domains.
Infeasibility identification for ADMM-based methods was a major concern, as it was established that the iterates of ADMM diverge if the problem is infeasible (Eckstein and Bertsekas, 1992). SCS is an ADMM-based method for general conic programs that solves a homogeneous self-dual embedding to identify infeasibilities (O’Donoghue et al., 2016). The work by Banjac et al. (2019) showed that the differences of the ADMM iterates are convergent and can be used to construct certificates of infeasibility for the original problem. This technique is used by both OSQP (Stellato et al., 2020) and COSMO (Garstka et al., 2021) to identify infeasible problems.
Unlike these approaches, FlexQP certifies infeasibility through its relaxation rather than as a separate detection mechanism, returning a minimum-violation solution directly. This addresses a weakness of the existing methods, as even though they can successfully identify infeasible problems, they do not return a useful approximate solution to the original problem that can be used by the user.
Sequential Quadratic Programming. Since its inception in 1963 (Wilson, 1963), SQP has evolved into a general methodology for solving complex and highly constrained NLP problems. A key challenge in SQP methods is that the QP subproblems formed by linearizing the constraints can become infeasible, even when the original NLP is feasible. This issue is particularly relevant for learning-based settings, where the iterative training procedure often causes intermediate QP subproblems to become infeasible.
Several strategies have been developed to handle such infeasibilities; a recent unified perspective of these methods is provided by Kiessling et al. (2025). One approach is to run a feasibility restoration phase to locate a feasible point when an infeasible QP is encountered (Nocedal and Wright, 2006), such as in the FilterSQP method by Fletcher and Leyffer (2002). However, these procedures require extensive tuning and are not scalable for batched settings. A complementary approach is to modify the QP subproblem itself to ensure that a meaningful step can always be computed. Burke and Han (1989) proposed a robust QP subproblem for this purpose, and SNOPT’s elastic mode (Gill et al., 2005) famously solves a relaxed QP when the original subproblem is infeasible. Meanwhile, stabilized SQP methods (Wright, 1998; Hager, 1999; Izmailov and Solodov, 2012) construct modified QP subproblems using an augmented Lagrangian objective, providing a mechanism to handle the degeneracy of practical optimization.
At the 2011 Advances in Numerical Computation Workshop, Gill (2011) noted that “almost all practical optimization problems are degenerate,” emphasizing the role that a robust QP solver plays in the field of NLP. Rather than handling infeasibilities and degeneracies on a per-problem basis, FlexQP unifies their treatment within a single batched solver, making it well-suited for learning-based SQP methods and GPU implementations. These considerations are particularly relevant for nonlinear MPC (Diehl et al., 2009; Rawlings et al., 2020) and safety-critical control (Ames et al., 2019; Wabersich and Zeilinger, 2021), where SQP is used in real-time control loops. A recent overview of constrained second-order trajectory optimization methods, including SQP and augmented Lagrangian-based approaches, is provided by Aoyama et al. (2024).
Learning to Optimize and Deep Unfolding. Learning to optimize or amortized optimization has established itself as a key technique for improving decision making using data-driven techniques (Chen et al., 2022b; Amos and others, 2023). Recent work has focused on understanding the representations (Liu et al., 2023), training dynamics (Metz et al., 2019; Scieur et al., 2022), and convergence (Sucker and Ochs, 2025) of learning-to-optimize approaches. These methods can be broadly classified into two main categories: model-free approaches that learn an optimizer from scratch, and model-based approaches that augment an existing algorithm with learnable components (Shlezinger and Eldar, 2023).
Deep unfolding, also known as algorithm unrolling (Monga et al., 2021), is a model-based learning-to-optimize approach that began with the seminal work by Gregor and LeCun (2010) on the learned iterative shrinkage and thresholding algorithm (LISTA). The LISTA framework has been extensively analyzed and improved over the years (Moreau and Bruna, 2017; Chen et al., 2018; Liu et al., 2019; Chen et al., 2021), and deep unfolding has since been applied to various fields, including speech processing (Hershey et al., 2014; Heigold et al., 2016), compressive imaging (Zhang and Ghanem, 2018; Mardani et al., 2018), and video reconstruction (Luong et al., 2021; De Weerdt et al., 2024), among others (Adler and Öktem, 2018; Gupta et al., 2018; Solomon et al., 2020). Deep unfolding has also been effectively applied to ADMM-based problems and algorithms (Yang et al., 2016; Ding et al., 2018; Xie et al., 2019; Yang et al., 2020; Saravanos et al., 2025).
Deep FlexQP is a deep-unfolded optimizer that learns hyperparameter policies from example problems of interest but maintains the mathematical structure from the underlying optimization process to inherit the guarantees and interpretability of the base optimizer.
Learned Solvers for Quadratic Programming. Learning-based approaches for quadratic programming have received considerable attention in recent years. Methods based on warm-starting determine a good initialization for a downstream solver using machine learning. Sambharya et al. (2023) learn a warm-start network for a Douglas-Rachford splitting QP solver, and the method was extended to general fixed-point algorithms in Sambharya et al. (2024). A related approach uses a deep-unfolded Douglas-Rachford splitting solver to warm-start SCS (Xiong et al., 2025).
A complementary class of approaches enhances the iterations of an optimization algorithm with learnable components. Deep FlexQP falls within this latter category of methods. Ichnowski et al. (2021) use reinforcement learning to learn dimension-agnostic policies for selecting the penalty parameters of OSQP. Saravanos et al. (2025) presented deep-unfolded optimization architectures for centralized and distributed QP, drawing an analogy to closed-loop control in order to design learned feedback policies for the hyperparameters of centralized and distributed OSQP. These approaches demonstrated remarkable scalability, solving problems with tens or even hundreds of thousands of variables to satisfactory accuracy under limited computational budgets. Similarly, Sambharya and Stellato (2024) learn step-varying and steady-state hyperparameters for several methods, including OSQP and SCS. We use the insight of dimension-independence from Ichnowski et al. (2021) along with the closed-loop control analogy from Saravanos et al. (2025) to learn component-wise LSTM policies for the hyperparameters of FlexQP. Consistent with the findings of Sambharya and Stellato (2024), the learned policies quickly adjust hyperparameters within the first few iterations before converging to a steady-state regime (see Appendix L).
Performance Guarantees for Learned Optimizers. Providing formal performance guarantees for learned optimizers remains an open challenge. Some asymptotic convergence guarantees have been derived using safeguarding mechanisms (Heaton et al., 2023) or by ensuring that the iterates do not deviate very far from a known convergent algorithm (Banert et al., 2024; Martin et al., 2025; Martin and Belgioioso, 2026). Other approaches ensure convergence through greedy objective descent (Fahy et al., 2024) or provide deterministic worst-case certificates via the performance estimation problem (PEP) framework (Sambharya et al., 2025).
Another class of guarantees relies on PAC-Bayes bounds (Sambharya and Stellato, 2025; Sucker et al., 2025; Saravanos et al., 2025; Sucker and Ochs, 2025), which provide certificates of performance that hold with high probability. However, the loss functions used in these works are based on binary objectives or linear ratios that saturate as the optimizer converges, yielding uninformative bounds when the learned optimizer achieves very low error. To address this, we develop our PAC-Bayes bounds based on a log-scaled, residual-based loss function that better captures the performance of the optimizer as the error gets arbitrarily small (see Section 4.3 and Appendix M).
Appendix B Proof of Theorem 3.1
@@ Appendix A Extended Related Work
Lemma B.1.
The relaxed QP Eq. 6 can equivalently be expressed as the following optimization:
$\operatorname*{minimize}_{x}\quad\phi(x;\mu_{I},\mu_{E})\vcentcolon=\frac{1}{2}x^{\top}Px+q^{\top}x+\mu_{I}\left\lVert(Gx-h)_{+}\right\rVert_{1}+\mu_{E}\left\lVert Ax-b\right\rVert_{1}.$ (15)
Proof.
Convert the optimization to the form without slack variables. This can be equivalently derived by directly relaxing the constraints of the original QP Eq. 3 using $\ell_{1}$ penalties. ∎
@@
    if $x^{*}$ solves Eq. 3, then $x^{*}$ solves Eq. 15, and
@@
    if $\hat{x}$ solves Eq. 15 then $\hat{x}$ solves Eq. 3.
@@
Let $x^{*}$ solve Eq. 3. For any $x\in\mathbb{R}^{n}$ we have that
@@
The step from Eq. 16b to Eq. 16c follows from the fact that $\mu_{I}\geq\left\lVert y_{I}^{*}\right\rVert_{\infty}$ and $\mu_{E}\geq\left\lVert y_{E}^{*}\right\rVert_{\infty}$, and Eq. 16d follows from the definition of the infinity norm. Eq. 16e follows from the definition of ReLU and the absolute value and Eq. 16f is implied by the linearity of the constraints. Obtaining Eq. 16g follows from the complementary slackness condition $y_{I,i}^{*}(g_{i}^{\top}x^{*}-h_{i})=0$ for all $i=1,\ldots,m$, and feasibility $a_{i}^{\top}x^{*}-b_{i}=0$ for all $i=1,\ldots,p$. Eq. 16i comes from the stationary condition $Px^{*}+q+G^{\top}y_{I}^{*}+A^{\top}y_{E}^{*}=0$. Finally, obtaining Eq. 16k follows from the positive semidefiniteness of $P$.
Thus, we have shown that $\phi(x^{*};\mu_{I},\mu_{E})\leq\phi(x;\mu_{I},\mu_{E})$ for any $x$, which implies that $x^{*}$ minimizes $\phi(x;\mu_{I},\mu_{E})$ and therefore solves Eq. 15.